# Extension of the integrated HydroKinetic Model to BES RHIC and GSI-FAIR nuclear collision energies

### v2

@@ II.3 Hydrodynamic expansion
The anticipated outcome of the hydrodynamic stage is the formation of a hypersurface that marks the transition between the fluid and gas phases, often referred to as the ’particlization hypersurface’, and denoted as $\sigma_{sw}$ in this paper. As we wrote before, we carry out this transition at a fixed energy density $\epsilon_{sw}$ for simplicity. In our model, the construction of the particlization hypersurface is achieved using the Cornelius routine [36, 37].
When dealing with low-energy collisions at $\sqrt{s_{NN}}\sim 3-10$ GeV, as observed in experiments such as HADES, RHIC BES, or CBM, it is reasonable to anticipate that the system may not achieve complete thermalization. Consequently, extending the particlization hypersurface criteria $\sigma_{sw}$ into the thermalization phase, $\tau<\tau_{th}$, becomes imperative.
@@
A few technical details. Minor time fluctuations can emerge in UrQMD tensors due to the transition from Cartesian to Milne coordinates during the thermalization stage, as in Eq. (13). They disappear if we average over several time steps. Consequently, as Eq. (18) implies, a smooth transition from the source to hydrodynamic tensors occurs. After several time steps, the hydrodynamic component behaves regularly over time, providing a smooth particlization hypersurface.
@@ II.4 Particlization
Following the transition through the particlization hypersurface, the matter becomes sufficiently dilute to be effectively described as a hadron-resonance gas. In the iHKMe, the injection of hadrons occurs from two distinct sources: equilibrium (eq) and non-equilibrium (n.eq)
$N_{\text{total}}=N_{\text{eq}}+N_{\text{n.eq}}.$ (25)
The particles from both components are injected into the afterburner cascade locally on the particlization hypersurface $\sigma_{sw}$ from the corresponding distribution function.
@@ II.4.1 Emission of thermal particles
The emission from the near-equilibrium source can be characterized using the well-established Cooper-Frye procedure [38] with small modifications discussed later. In this prescription, the yield of hadrons of the species $i$ is given by:
$N^{i}_{\text{eq}}=\int d^{3}p\int\frac{d\sigma_{\mu}p^{\mu}}{p_{0}}\left(1-{\cal P}(\tau)\right)f_{i}(x,p),$ (26)
where $d\sigma_{\mu}$ is a small part of the hypersurface $\sigma_{sw}$, and $f_{i}$ is the near-equilibrium distribution function. Notice here the additional $1-{\cal P}(\tau)$ factor, which is not equal to one during the relaxation stage.
The distribution function $f_{i}(x,p)$ depends on the local thermodynamic properties of the system, temperature, chemical potential, and shear stress tensor. The velocity of the fluid $u_{\mu}$, the shear-stress tensor $\pi^{\mu\nu}(x)$, and the baryon charge density $n_{B}$ are extracted from the hydrodynamics energy-momentum tensor and baryon current. The baryon, electric, and strange chemical potentials, as well as the temperature, are obtained from the energy density $\epsilon$ and the charge densities $n_{b}$, $n_{q}=0.4n_{b}$, $n_{s}=0$, using the equation of state. The chemical potential $\mu_{i}$ for each particle species can then be determined from the following relation:
@@
where $B_{i}$, $q_{i}$, and $s_{i}$ represent the particle’s baryon, electric, and strange charges, respectively. The near-equilibrium distribution functions $f_{i}(x,p)$ are obtained after applying the Grad ansatz [39] for viscous corrections to the corresponding local equilibrium distribution functions $f^{eq}_{i}(x,p)$. Assuming the same corrections for all hadron species, the thermal particle production can be written as:
@@
×$\quad\times\left(1+(1\mp f_{\text{eq}})\frac{p_{\mu}p_{\nu}\pi^{*\mu\nu}}{2T^{2}(\epsilon+p)}\right),$
Here, the asterisk denotes the local rest frame of the fluid element, and $\mp$ indicates Fermi/Bose statistics. For more detailed information, we refer the reader to the papers [12] or [40].
@@ II.4.2 Non-thermal emission
The non-equilibrium component is emitted similarly.
$N^{i}_{\text{n.eq}}=\int d^{3}p\int\frac{d\sigma_{\mu}p^{\mu}}{p_{0}}{\cal P}(\tau)f^{\text{n.eq}}_{i}(x,p).$ (29)
However, instead of generating particles from the distribution function $f^{n.eq}_{i}(x,p)$ we utilize the particles from the sub-ensemble constructed during the pre-equilibrium dynamics stage regarding non-thermal emission. Then, for each piece of the particlization hypersurface $d\sigma^{\mu}_{j}$ located at $x^{\mu}_{j}$, the probability that a particle with space-time position $x^{\mu}_{i}$ will be emitted is defined by the product:
${\cal G}(p_{i},x_{i};d\sigma_{j})$ $=\theta(p_{i}^{\mu}d\sigma^{j}_{\mu})\times$ (30)
$\theta\left(\frac{\Delta\tau}{2}-|\tau_{j}-\tau_{i}|\right)\times{\cal K}_{ij}$$\times\frac{p_{i}^{\mu}d\sigma^{j}_{\mu}}{p^{0}_{i}}.$
The first Heaviside step ensures that particles fly from the hotter to the colder phase. The following two multipliers come from the construction of the nonequilibrium distribution function in Eqs. (8) and (13). These terms determine whether the particle intersects the hypersurface in space and time. The last term is related to the size and space-time orientation of the hypersurface. Summing over the whole hypersurface and all particle tracks with the ${\cal P}_{\tau}$ weight yields the total contribution of non-thermal emission.
$N_{\text{n.eq}}=\sum_{i\in\{tracks\}}\sum_{j\in\{\sigma_{sw}\}}{\cal P}_{\tau_{j}}{\cal G}(p_{i},x_{i};d\sigma_{j}).$ (31)
@@ II.5 Hadronic cascade
At the final post-hydrodynamical stage of the system’s evolution, all particles from both equilibrium and non-equilibrium sources are input into the UrQMD hadron cascade code [26]. In the iHKM framework, we aim to account for all reliably known hadron resonance states, even those not processed by UrQMD. Therefore, heavy resonances not present in the UrQMD particle database are decayed right at $\sigma_{sw}$ to ensure energy-momentum conservation.
We generate 20 to 200 UrQMD events based on a single hydrodynamic run to increase statistics in event-by-event simulation. A detailed discussion of this procedure can be found in [43]. It’s worth noting that this approach saves CPU time while, as demonstrated in similar models such as [40] and the iHKM analysis, it does not significantly impact the final observables, including Bose-Einstein correlations. However, we expect that artificial correlations might be present at several GeV energies when the multiplicities of thermal particles are relatively low due to this procedure. Therefore, it is reasonable to stick with a pure event-by-event simulation. However, this should not be the case for the $14.5$ GeV energies considered in all the simulation results presented in this paper.
@@ III.1 Smoothing procedure
Figure 1: Energy density distribution in the transverse plane at $z=0$ (i.e., zero space rapidity $\eta=0$). The same event of Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV from the 20-30% centrality class is shown. Left: energy density extracted from the UrQMD tensor (14) at $\tau=\tau_{th}=2.6$ fm/$c$. Right: energy density in iHKM simulations at the same proper time $\tau_{th}$.
@@ III.2 Free parameters and calibration
    • $\tau_{0}$ start of thermalization stage
    • $\tau_{rel}$ relaxation time
    • $\tau_{th}$ end of thermalization stage
@@
    • particlization energy density $\epsilon_{sw}$
@@
For the transition to the afterburner stage, we use a typical value for the energy density $\epsilon_{sw}=0.5$ GeV/$c$, which lies in the range where the equations of state for the liquid and gas phases of strongly interacting matter coincide. However, reducing this value might enhance the results due to the less intense hadron annihilation process in UrQMD compared to local-equilibrium hydrodynamics. Specifically, we observe the sensitivity of $\bar{p}/p$ to $\epsilon_{sw}$, prompting us to treat $\epsilon_{sw}$ as a free parameter if necessary. The final sets of parameters we employed to simulate Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV with two different equations of state are presented in Table 1.

### v3

@@ II.3 Hydrodynamic expansion
The anticipated outcome of the hydrodynamic stage is the formation of a hypersurface that marks the transition between the fluid and gas phases, often referred to as the ’particlization hypersurface’, and denoted as $\sigma_{\text{sw}}$ in this paper. As we wrote before, we carry out this transition at a fixed energy density $\epsilon_{\text{sw}}$ for simplicity. In our model, the construction of the particlization hypersurface is achieved using the Cornelius routine [50, 51].
When dealing with low-energy collisions at $\sqrt{s_{NN}}\sim 3-10$ GeV, as observed in experiments such as HADES, RHIC BES, or CBM, it is reasonable to anticipate that the system may not achieve complete thermalization. Consequently, extending the particlization hypersurface criteria $\sigma_{\text{sw}}$ into the thermalization phase, $\tau<\tau_{th}$, becomes imperative.
@@
The transport model UrQMD employs an EoS corresponding to a hadron resonance gas, whereas the hydrodynamic stage may use a different EoS, such as one incorporating a QCD crossover or a phase transition. This mismatch in EoS can lead to so-called matching artifacts, including discontinuities in pressure, entropy non-conservation, and spurious flows at the transition hypersurfaces. While energy and momentum conservation can be ensured during the mapping procedures, the differences in thermodynamic assumptions between models can introduce systematic effects in the final-state observables. These artifacts are a known limitation of hybrid approaches and should be kept in mind when interpreting results.
@@
A few technical details are worth noting. Minor time fluctuations can appear in the UrQMD energy-momentum tensor components due to the transformation from Cartesian to Milne coordinates during the thermalization stage, as described in Eq. (13) and discussed further in Appendix A. These fluctuations vanish when averaging over several time steps. As a result, the transition from the particle-based source to the hydrodynamic energy-momentum tensor, governed by Eq. (18), becomes smooth. Empirically, we find that averaging over eight consecutive time steps is sufficient to ensure that the hydrodynamic component evolves smoothly in time, ultimately leading to a well-behaved particlization hypersurface.
It is also important to note that at the initial proper time when the construction of the particlization hypersurface begins, some regions of the system already have an energy density below the particlization threshold $\epsilon_{\text{sw}}$. These regions are typically located at the system’s periphery (the so-called corona) or in the spectator zones. Particles in these areas generally do not undergo thermalization and moreover, are ignored by the Cornelius module.
To ensure the conservation of total energy, momentum, and conserved charges in the system, we manually extend the particlization hypersurface to include these non-thermalized regions. For such extensions, we assign hypersurface elements a normal vector corresponding to constant proper time, as given in Eq. (9), and set $\mathcal{P}_{\tau}=1$ for them—indicating that these regions are purely non-equilibrated.
@@ II.4 Particlization
Following the transition through the particlization hypersurface, the matter becomes sufficiently dilute to be effectively described as a hadron-resonance gas. In the current version of iHKM, hadrons are injected from two distinct sources: the equilibrium (eq) and non-equilibrium (n.eq) components. Analogous to the decomposition in Eqs. (1415), the system’s distribution function is expressed as the sum of these two contributions:
$f(x,p)=f_{\text{n.eq}}(x,p)\cdot\mathcal{P}_{\tau}+f_{\text{eq}}(x,p)\cdot\left(1-\mathcal{P}_{\tau}\right)\,$, (25)
where $f_{\text{n.eq}}(x,p)$ corresponds to non-equilibrated hadrons originating from the same UrQMD event used for the initial dynamics and is constructed using the kernel (8), while $f_{\text{eq}}(x,p)$ represents the standard locally near-equilibrium component typically used in hydrodynamic-to-transport switching procedures. Notably, once the thermalization time is reached, $\mathcal{P}_{\tau}\to 0$, and all particles are emitted exclusively from the near-equilibrium source, consistent with conventional hybrid models.
Both components of the distribution function can be expressed as a sum over all hadron species $i$:
$f_{\text{eq}/\text{n.eq}}(x,p)=\sum_{i}f^{\text{eq}/\text{n.eq}}_{i}(x,p)\,$, (26)
where, for the equilibrium component, the sum runs over all hadron species in the ideal hadron-resonance gas into which the hydrodynamic fluid is converted. For the non-equilibrium component, the sum includes all hadron species present in the UrQMD event used to generate the initial conditions.
For each infinitesimal element of the particlization hypersurface $\Delta\sigma^{\mu}(x)$, particles are then generated using the well-established Cooper–Frye prescription [52], with minor modifications discussed later:
$N_{i}(x,p)=\frac{\Delta\sigma_{\mu}p^{\mu}}{p^{0}}f_{i}(x,p)\,.$ (27)
Particles from both the equilibrium and non-equilibrium components are sampled independently using their respective distribution functions, $f_{\text{eq}}$ and $f_{\text{n.eq}}$, without applying the factors $\mathcal{P}_{\tau}$ and $1-\mathcal{P}_{\tau}$ directly during sampling. These weights are instead used to determine the relative fraction of particles from each component that are subsequently passed into the UrQMD afterburner stage.
@@ II.4.1 Emission of thermal particles
The near-equilibrium distribution function $f^{\text{eq}}_{i}(x,p)$ for each particle species depends on the local thermodynamic properties of the system: temperature, chemical potentials, and the shear-stress tensor. The fluid velocity $u^{\mu}(x)$, energy density $\epsilon$, and baryon charge density $n_{B}(x)$ are extracted from the hydrodynamic energy-momentum tensor, the net baryon current, and the equation of state using the standard Landau matching procedure:
$T^{0\mu}_{\text{hydro}}$ $=(\epsilon+p)\,u^{0}u^{\mu}-p\,g^{0\mu},$ (28)
$J_{\text{hydro}}^{0}$ $=n_{B}u^{0},$
$p$ $=p(\epsilon,n_{B})\,.$
Note that these tensors are obtained from the tilded ones—derived from the hydrodynamic equations with source terms, Eqs. (18)–(19)—by dividing them by the factor $1-{\cal P}_{\tau}$, as shown in Eqs. (23) and (24). Also, the components $\pi^{0\mu}$ vanish here because the shear-stress tensor is orthogonal to the fluid four-velocity in the Landau frame [47].
However, when transitioning from the hydrodynamic description to the hadron gas phase, the corresponding equations of state (EoS) typically do not match exactly. This mismatch results in discontinuities in the thermodynamic properties of the system across the particlization hypersurface. In our model, we adopt the hadron gas EoS to solve Eqs. (28) and use the corresponding energy density as the switching criterion within the Cornelius routine. This approach ensures local—and therefore global—conservation of all conserved charges. However, it introduces small discontinuities: typically less than 1% in $\epsilon$ and $n_{B}$, and around 10% in pp. [^2] In a statistically insignificant number of cases, these discontinuities may be much larger due to failures of the numerical algorithms.
On the other hand, closeness of intensive quantities leads to noticeable discontinuities in extensive properties such as temperature $T$ and baryon chemical potential $\mu_{B}$. This situation can be improved by employing an equation of state for the hydrodynamic stage that smoothly connects to the hadron resonance gas properties across a wide range of the QCD phase diagram. In future work, we plan to implement one such EoS (NEOS BQS [53]) into our model.
When the local temperature and the baryon, electric, and strange chemical potentials are determined from the local energy density $\epsilon$ and the conserved charge densities $n_{B}$, $n_{Q}=0.4\,n_{B}$, and $n_{S}=0$ using the equation of state, one can construct the equilibrium distribution function for the ideal hadron resonance gas. The chemical potential $\mu_{i}$ for each particle species is then given by:
@@
where $B_{i}$, $q_{i}$, and $s_{i}$ represent the particle’s baryon, electric, and strange charges, respectively. The near-equilibrium corrections to the distribution functions are obtained after applying the Grad ansatz [54] for viscous corrections. Assuming the same corrections for all hadron species, the thermal particle production in the rest frame of the fluid can be written as:
@@
$\quad\times\left(1+(1\mp f_{\text{eq}})\frac{p_{\mu}p_{\nu}\pi^{\mu\nu}}{2T^{2}(\epsilon+p)}\right),$
Here, $\mp$ indicates Fermi/Bose statistics. For more detailed information, we refer the reader to the papers [12] or [55].
@@ II.4.2 Non-thermal emission
As mentioned above, for the non-equilibrium distribution function, we use the same set of particles from the sub-ensemble constructed during the pre-equilibrium dynamics stage. Then, for each element of the particlization hypersurface $d\sigma_{j}^{\mu}$ located at position $x_{j}^{\mu}$, the probability that a particle with space-time coordinate $x_{i}^{\mu}$ will be emitted is given by the product:
${\cal W}(p_{i},x_{i};d\sigma_{j})$ $=\theta\bigl(p_{i}^{\mu}d\sigma^{j}_{\mu}\bigr)\times\theta\left(\frac{\Delta\tau}{2}-\left|\tau_{j}-\tau_{i}\right|\right)$ (31)
$\quad\times{\cal K}_{ij}\times\frac{p_{i}^{\mu}d\sigma^{j}_{\mu}}{p_{i}^{0}}.$
The first Heaviside step function ensures that particles move from the hotter to the colder phase. The next two factors arise from the construction of the non-equilibrium distribution function in Eqs. (8) and (13), determining whether the particle intersects the hypersurface in space and time. The last term accounts for the size and space-time orientation of the hypersurface element. Summing over the entire hypersurface and all particle trajectories from the initial UrQMD event, weighted by ${\cal P}_{\tau}$, yields the total contribution of non-thermal emission:
$N_{\text{n.eq}}=\sum_{i\in\{\text{tracks}\}}\sum_{j\in\{\sigma_{\text{sw}}\}}{\cal P}_{\tau_{j}}\,{\cal W}(p_{i},x_{i};d\sigma_{j}).$ (32)
@@ II.4.3 Non-space-like surface emission treatment
II.4.4 Fluidization and Particlization Rates
Having discussed the mechanisms of converting particles to fluid (thermalization or fluidization) and the reverse process (particlization), we now turn to examining the rates of these processes in typical simulations. To describe these qualitatively, we calculate the ratio of the energy stored in the hydrodynamic part of the system—corresponding to the stress-energy tensor given by Eq. (20) (including the factor $\mathcal{P}_{\tau}$)—to the total energy of the system, Eq. (14). The energy (in the laboratory frame) is computed as the flux of the $T^{0\mu}$ component of the stress-energy tensor through the hypersurface of constant proper time.
$E(\tau)=\int_{\tau}d\sigma_{\mu}\,T^{0\mu}\,$, (35)
where for the hydrodynamic component we additionally exclude parts where $\epsilon<\epsilon_{\text{sw}}$ as they are needed for energy conservation and smooth boundary conditions for hydrodynamics, but they correspond to already hadronized matter. The ratio $E_{\text{hydro}}/E_{\text{total}}$ as a function of $\tau$ is shown in Fig. 1. The general trend is as follows:
  • Before the onset of the thermalization stage ($\tau<\tau_{0}$), the entire system is out of equilibrium and is described by the UrQMD cascade. In this work, we neglect the slow, “natural” thermalization that occurs within UrQMD (see [58]).
  • During the relaxation stage ($\tau_{0}<\tau<\tau_{\text{th}}$), the dense regions of the system begin to thermalize, leading to a gradual increase in the hydrodynamic component.
  • At the periphery of the system, fluid elements begin to hadronize, reducing the overall hydrodynamic fraction. As a result, the maximum contribution of the hydrodynamic component typically occurs before $\tau_{\text{th}}$. Figure 1 shows that the maximum fluid fraction depends on the details of the thermalization process—namely, the values of $\tau_{0}$ and $\tau_{\text{th}}$—as well as on the collision energy $\sqrt{s_{NN}}$. Earlier thermalization and higher collision energies lead to a larger portion of the system reaching local equilibrium.
  • After the thermalization time $\tau_{\text{th}}$—which may not be reached at very low collision energies or in non-central collisions—the system continues to expand, cool, and hadronize until the maximum energy density drops below the switching threshold $\epsilon_{\text{sw}}$.
  • Beyond this point, the entire system is described as a hadron gas.
In this context, let us briefly discuss the similarities and differences between our approach and other hybrid models. Similar to our model, the frameworks presented in [59] and [55] employ transport models (SMASH and UrQMD, respectively) and operate in Milne coordinates. However, unlike our approach, these models do not include an explicit thermalization stage. Instead, they assume instantaneous thermalization, typically occurring at the time of nuclear overlap (see Eq. (36) in the next chapter). In our model we expect hierarchy of times $\tau_{0}<\tau_{\text{overlap}}<\tau_{\text{th}}$. The differences between the models become increasingly significant at lower collision energies, particularly below 5 GeV, where we expect only partial thermalization of the system. Moreover, dissipative effects during the early stages are expected to be more pronounced in our model due to the non-equilibrium component, which may lead to different inferred values for transport coefficients.
Compared to the version of iHKM developed for ultrarelativistic energies [15], which leads to full thermalization of the bulk of the system at some hypersurface, the current version additionally incorporates a smooth particlization of the system during this stage. At LHC energies, the matter density is significantly higher and thermalization occurs much earlier; as a result, the effects of gradual particlization were negligible due to their small magnitude.
In contrast to other models with continuous fluidization [37, 38, 36], we do not explicitly prescribe a specific mechanism for thermalization. As a result, our model lacks predictive power in certain aspects; however, it is more general and flexible. In principle, our approach allows for the possibility of constraining the thermalization mechanism by implementing more sophisticated forms of the function $\cal P$} (see Eq. 14), for example, by allowing it to depend on local energy density. Such extensions, however, have not yet been implemented.
Concerning the thermalization rate, Ref. [37] presents the evolution of the fluid fraction as a function of time at various collision energies, which enables a direct comparison with our model. It appears that at lower energies (below $\sqrt{s_{NN}}\sim 10$ GeV), the maximum fluid fraction is reached near the nuclei overlap time, while at higher energies it is attained later. This behavior is particularly important for the development of hydrodynamic flow and therefore significantly affects key observables such as transverse momentum spectra, flow harmonics, and the sizes of the particle-emitting source extracted via interferometric analysis. In particular, for $\sqrt{s_{NN}}=14.5$ GeV—the energy used for demonstration in this work—the hydrodynamic phase in central collisions is expected to end around 99 fm/$c$, while in Ref. [37], this duration can exceed $15$ fm/$c$ in the central region of the system.
Another notable difference from other models concerns the treatment of core–corona separation, which is implemented in the aforementioned works as well as in other studies (e.g., [60]). In iHKM, the corona is also spatially and temporally separated at the periphery of the system, similar to those models. However, a non-thermal component is also present in the dense region of the system, in the sense of ensemble averaging.
Figure 1: Ratio of the energy stored in the hydrodynamic part of the system to the total energy as a function of $\tau$. For each of the lines, one event from 050-5 % centrality is considered.
@@ II.5 Hadronic cascade
At the final post-hydrodynamical stage of the system’s evolution, all particles from both equilibrium and non-equilibrium sources are input into the UrQMD hadron cascade code [26]. In the iHKM framework, we aim to account for all reliably known hadron resonance states, even those not processed by UrQMD. Therefore, heavy resonances not present in the UrQMD particle database are decayed right at $\sigma_{\text{sw}}$ to ensure energy-momentum conservation.
We generate 20 to 200 UrQMD events based on a single hydrodynamic run to increase statistics in event-by-event simulation. A detailed discussion of this procedure can be found in [61]. It’s worth noting that this approach saves CPU time while, as demonstrated in similar models such as [55] and the iHKM analysis, it does not significantly impact the final observables, including Bose-Einstein correlations. However, we expect that artificial correlations might be present at several GeV energies when the multiplicities of thermal particles are relatively low – due to this procedure. Therefore, it is reasonable to stick with a pure event-by-event simulation. However, this should not be the case for the $14.5$ GeV energies considered in all the simulation results presented in this paper.
@@ III.1 Smoothing procedure
Figure 2: Transverse-plane distribution of the energy density (logarithmic scale) at $z=0$ for a single Au+Au collision event at $\sqrt{s_{NN}}=14.5$ GeV in the 20–30% centrality class. Left: energy density extracted from the UrQMD tensor (11), constructed using a Gaussian kernel (8) with $R=0.5$ fm, at the thermalization proper time $\tau=\tau_{\rm th}=2.6$ fm/$c$. Right: energy density obtained from the hydrodynamic tensor in iHKM simulations at the same $\tau_{\rm th}$ (other model parameters chosen according to Set 1 in Table 1). Black dashed lines separate regions with energy density above and below $0.5$ GeV/fm3.
@@ III.2 Free parameters and calibration
    • $\tau_{0}$ - start of thermalization stage
    • $\tau_{rel}$ - relaxation time
    • $\tau_{th}$ - end of thermalization stage
@@
    • particlization energy density $\epsilon_{\text{sw}}$
@@
For the transition to the afterburner stage, we use a typical value for the switching energy density, $\epsilon_{\text{sw}}=0.5$ GeV/fm3, which lies in the range where the equations of state (EoS) for the liquid and gas phases of strongly interacting matter are close. However, reducing this value may improve the results, due to the less intense hadron annihilation processes in UrQMD compared to local-equilibrium hydrodynamics. In particular, we observe a noticeable sensitivity of the $\bar{p}/p$ ratio to $\epsilon_{\text{sw}}$, which motivates us to treat $\epsilon_{\text{sw}}$ as a free parameter when necessary. Similar sensitivity of this ratio to the details of the EoS and the particlization energy density has been reported in other studies (see, e.g., [53]), and has been addressed in numerous theoretical works [65, 66]. The final sets of parameters used to simulate Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV with two different equations of state are summarized in Table 1.